OpenStudy (anonymous):

formula for last 2 questions in H8P3??? can any1 help me pls??

OpenStudy (anonymous):

answers for last 2 questions-52 34

OpenStudy (anonymous):

don't work

OpenStudy (anonymous):

it didnt work out....can u lemme kno d formulas

OpenStudy (anonymous):

don't work for me too need the formulae

OpenStudy (anonymous):

to all above users manipulate the values subtracting 1 or adding 1 to eACH VALUE

OpenStudy (anonymous):

what do u mean???

OpenStudy (anonymous):

what do u mean??? [2]

OpenStudy (anonymous):

I didnt get u @aditya07

OpenStudy (anonymous):

The last two answers are 74 and 29.

OpenStudy (anonymous):

any of the above repsonses not working out ..!can any one tell me the formulas

OpenStudy (anonymous):

the last one is 29.5159549597

OpenStudy (anonymous):

try this one, it works for me...

OpenStudy (anonymous):

its nt workin

OpenStudy (anonymous):

3RD---(-ln( (3.5-5) / (0.5-5) ) * (RPU+RON) * 3.5*10^-15) * 10^12 4TH----(-ln( (0.9-(5*1800/(RON+RPU))) / (3.5-(5*RON/(RON+RPU))) ) * ((RON|| RPU) + RON) * 3.5*10^-15) * 10^12

OpenStudy (anonymous):

what is the last ans. of H8P3: MEMORY

OpenStudy (anonymous):

H8P1 IMPULSE (a) What is the value of vR1(t) at t=0− in Volts (V)? (b) What is the value of vR1(t) at t=0+ in volts (V)? Hint: Recall that that ∫0+0−δ(t)dt=1 (c) What is the value for vR1(t) at t=1ms in volts(V)? (Hint: The L-R1 and C-R2 branches of the circuit are decoupled.) ? ? ?

OpenStudy (anonymous):

62 and 35 worked for me

OpenStudy (anonymous):

use 24 its worked for me if it will worked for u then give me best response plzzzzz